it’s linear control system

R-YS T Nay1 Plant ontrol o J.9 S(St29) TN Lind s)
Answer #1

Plant P(s)ss+20

Controller C(s) = K

Feedback Y(s) N(s)

-(Y(s)N(s) E

E* C(s)P(s) Y(s)

Substituting E(s):

(Y(s)N(s)) C(s) * P(s) Y(s)

-YCP-NCP = Y

NCP YYCP

NCP Y(1+CP)

-CP N (1CP)

substituting C(s) and P(s)

\frac{Y}{N}=\frac{-K \frac{1}{s(s+20)}}{(1+K \frac{1}{s(s+20)})}

\frac{Y(s)}{N(s)}=\frac{-K }{s(s+20)+K}